# Torsion points on $\text{GL}_{2}$-type abelian varieties

### v1

@@ 1. Introduction
In this paper, we work more generally on abelian varieties of $\text{GL}_{2}$-type defined over a number field $K$, without further restrictions on the dimension and on the endomorphisms. Our strategy is to use a classical approach in the study of rational points, that is, answering to a local-global problem. Indeed, for any abelian variety $A/K$, if $\mathfrak{p}$ is a prime of good reduction for $A$, the group $\operatorname{Tors}(A)(K)$ injects into the finite group of $\mathbb{F}_{\mathfrak{p}}$-points of the reduction of $A$ modulo $\mathfrak{p}$. Thus we have that $|\operatorname{Tors}(A)(K)|$ divides $N(p):=|\tilde{A}(\mathbb{F}_{\mathfrak{p}})|$, where $\tilde{A}$ is the reduction of $A$ modulo $\mathfrak{p}$. A famous question by Lang asks if the converse is also true. More precisely he asked the following.
@@
Finally, we show how one can apply our main results to study the torsion of modular abelian varieties defined over the rationals, which, by Serre’s conjectures, correspond to weight $2$ newforms. We implement our results in the computational algebra system Magma, in order to produce tentative lists of possible torsion orders for modular abelian abelian varieties of dimension up to $5$, which are presented in Conjectures 4.4 and 4.5. The code and the outputs can be found at [AC26]. Although it is not currently known if the lists are exhaustive, we remark that all the torsion orders appearing in the lists occur for at least one abelian variety, which can be exhibited in terms of the corresponding newform. This data enhances the data that is currently available in LMFDB [17]. For example, there exists at least one abelian surface in the isogeny class associated to the level $39$ newform with LMFDB label 39.2.a.b which has torsion order equal to $28$. This abelian variety is not listed in LMFDB and there are no abelian surfaces with torsion order equal to $28$ in the database either.
@@ 2. Setting
Note that, a priori, $P_{\mathfrak{p}}^{\lambda}(t)$ has coefficients in $F_{\lambda}$, but since the representations $\rho_{\lambda}$ are $F$-rational, in fact $P_{\mathfrak{p}}^{\lambda}(t)\in F[t]$.
@@ 3. Main results
Let $A$ be an abelian variety over a number field $K$ of $\text{GL}_{2}$-type. Let $\ell$ be a rational prime such that, for all places of $K$ above $\ell$, $A$ has good reduction modulo such places (we simply say that $\ell$ is of good reduction for $A$). Let $\lambda$ be a prime of $F$ over $\ell$ and suppose that, for an integer $n\geq 1$, the congruence
@@
First, let $A^{\prime}$ be an abelian variety that is $K$-isogenous to $A$. By injectivity of the reduction map, which holds for all primes in $\Sigma$ (see [KAT81, Appendix]), $\operatorname{Tors}(A^{\prime})(K)$ injects into the $\mathbb{F}_{\mathfrak{p}}$-points of $A^{\prime}$ for all $\mathfrak{p}\in\Sigma$, thus $|\operatorname{Tors}(A^{\prime})(K)|$ divides $|A^{\prime}(\mathbb{F}_{\mathfrak{p}})|=N(\mathfrak{p})$. Taking the sup over all $A^{\prime}$ that are $K$-isogenous to $A$ and the gcd of all $N(\mathfrak{p})$ as $\mathfrak{p}$ varies in $\Sigma$, we obtain
$\sup_{A^{\prime}K\text{-isogenous to}A}|\operatorname{Tors}(A^{\prime})(K)|\big|\gcd_{\mathfrak{p}\in\Sigma}N(\mathfrak{p})$
Now, let $\ell$ be a prime of good reduction for $A$ dividing all $N(\mathfrak{p})$ for $\mathfrak{p}\in\Sigma$. As shown in Equation (1),
@@
The last equality follows from the compatibility of the representations $\{\rho_{\lambda}\}_{\lambda}$. Since this quantity is positive, there exists $\lambda_{0}$ lying over $\ell$ with $v_{\lambda_{0}}(P_{\mathfrak{p}}^{\lambda_{0}}(1))=n>0$. By Theorem 3.2, there exists an abelian variety $A^{\prime}$ that is $K$-isogenous to $A$ and such that $v_{\ell}(|\operatorname{Tors}(A^{\prime})(K)|)\geq n\mathfrak{e}_{\lambda_{0}}\mathfrak{f}_{\lambda_{0}}>0$, concluding the proof of the first statement.
@@
So, $v_{\ell}(\gcd N(\mathfrak{p}))\leq n\mathfrak{f}_{\lambda_{0}}\leq v_{\ell}(|\operatorname{Tors}(A^{\prime})(K)|)$.
@@ 4. Modular abelian varieties defined over $\mathbb{Q}$
We now specialise the result of Theorem 3.2 and Corollary 3.3 to $\text{GL}_{2}$-type abelian varieties defined over the rational numbers. Let $A$ be an abelian variety of dimension $g$ defined over $\mathbb{Q}$. Let $\operatorname{End}^{0}_{\mathbb{Q}}(A)=\operatorname{End}_{\mathbb{Q}}(A)\otimes_{\mathbb{Z}}\mathbb{Q}$ denote the algebra of endomorphisms of $A$ that are defined over $\mathbb{Q}$, and let $\operatorname{End}_{\overline{\mathbb{Q}}}^{0}(A)=\operatorname{End}_{\overline{\mathbb{Q}}}(A\times_{\mathbb{Q}}\overline{\mathbb{Q}})\otimes_{\mathbb{Z}}\mathbb{Q}$ denote the full endomorphism algebra.
@@
We now investigate the possible full endomorphisms algebras for simple abelian varieties of $\text{GL}_{2}$-type over $\mathbb{Q}$. As in §2, we write $F\hookrightarrow\operatorname{End}_{\mathbb{Q}}^{0}(A)$, where $F$ is a degree $g$ number field. By [RIB04, Theorem 2.1], if $A/\mathbb{Q}$ is simple, then $F=\operatorname{End}_{\mathbb{Q}}^{0}(A)$.
@@
Let us assume that $A$ is geometrically simple, so that $\operatorname{End}^{0}_{\overline{\mathbb{Q}}}(A)$ is an Albert algebra. If it is of type IV, then by Shimura’s decomposition [RIB04, §5], $A_{\overline{\mathbb{Q}}}$ is a power of a CM elliptic curve, and in particular it is not geometrically simple, contradiction. Then, by Albert’s classification, if it $A$ does not have PQM, then necessarily $\operatorname{End}^{0}_{\overline{\mathbb{Q}}}(A)$ is of type I, i.e. it is a totally real number field of degree dividing $g$. Then we have
$F\subseteq\operatorname{End}^{0}_{\mathbb{Q}}(A)\subseteq\operatorname{End}^{0}_{\overline{\mathbb{Q}}}(A),$
with $F$ as above. Thus all inclusions are equalities. ∎
By [RIB04, Theorem 4.4] and Serre’s modularity conjectures [SER87] (proved by Khare–Winterberger, [KW09a, KW09b]), a simple abelian variety $A/\mathbb{Q}$ of $\text{GL}_{2}$-type is modular, hence it is isogenous to the abelian variety $A_{f}$ associated to some newform $f$, with the notation used in [RIB04]. An explicit construction of $A_{f}$ can be found in [DS05, §9], as well as the definition of $\lambda$-adic representations attached to $f$ and to $A_{f}$. It is well known [DS05, Theorem 9.6.5] that the representations $(\rho_{f,\lambda})$ form a strictly compatible system for all $\lambda$’s, including those of bad reduction for $A_{f}$. Recall that $F=\operatorname{End}_{\mathbb{Q}}(A_{f})$ is equal to the coefficients field of $f$.
Therefore, for $A/\mathbb{Q}$ of $\text{GL}_{2}$-type, we can restate Theorem 3.2 and Corollary 3.3 as follows.
@@ 4.1. Implementation and conjectural results
In [AC26], we implement in Magma the results of Theorem 4.2 and Corollary 4.3 to produce a tentative list of torsion orders for modular abelian varieties of dimension $g$ over $\mathbb{Q}$. The idea is to loop over all newforms of weight $2$ and level up to a certain bound, consider the newforms having dimension (i.e. degree of the coefficient field over $\mathbb{Q}$) equal to deg =g=g, and compute, for each, the greatest common divisor of $P_{p}^{\lambda}(1)=\det(1-\rho_{f,\lambda}(\operatorname{Frob}_{p}))$, as $p$ varies among the primes of good reduction for $A$ between $2$ and a certain upper bound st, such that $\mathfrak{e}_{p}<p-1$. Ideally, we include enough primes so that the g.c.d. of the finitely many $P_{p}^{\lambda}(1)$’s we consider is equal to the actual g.c.d. computed using all primes. We suspect that st equal to the Sturm bound would suffice, and we have computational evidence to support this idea, but could not find a reference for that.
The code computes the predicted torsion order, i.e. the product
 $\prod_{\ell\text{ prime}}\ell^{\sup_{\lambda\mid\ell}(\mathfrak{f}_{\lambda}n)},\quad n\text{ maximal such that }P_{p}^{\lambda}(1)\equiv 0\pmod{\lambda^{n}}$,
@@
The predicted torsion order is not always equal to the greatest common divisor of the $N(p)$’s, hence Theorem 4.2 and Corollary 4.3 cannot be improved to be exactly equivalent to Katz’s result. However, up to conductor $500$, the two lists computed by the code are exactly the same for abelian varieties of dimension $2$ and $3$.
We conclude by stating two conjectures, which are suggested by the outcome of the code, on the possible torsion orders of abelian varieties of $\text{GL}_{2}$-type and dimension up to $3$, and on the primes dividing such torsion orders. All the numbers in the table are divisors of the predicted torsion orders computed by the code, and the list of all newforms with level up to $500$ and dimension up to $3$ and the respective predicted torsion orders can be found at [AC26]. More complete data on the predicted torsion orders for $g$ up to $5$ is available, and constantly updated. We do not currently know whether the lists in Conjectures 4.4 and 4.5 are exhaustive.
@@
Let $A/\mathbb{Q}$ be a $g$-dimensional abelian variety of $\text{GL}_{2}$-type. The possible orders of $\operatorname{Tors}(A)(\mathbb{Q})$ are listed below.
@@
Let $A/\mathbb{Q}$ be a $g$-dimensional abelian variety of $\text{GL}_{2}$-type and let $\ell$ be a prime such that $A$ admits a $\ell$-torsion point.
  • If $g=2$, then $\ell\in\{2,3,5,7,11,13,19\}$.
  • If $g=3$, then $\ell\in\{2,3,5,7,11,13,17,23,29,31\}$.

### v3

@@ 1. Introduction
In this paper, we work more generally on abelian varieties of $\text{GL}_{2}$-type defined over a number field $K$, without further restrictions on the dimension and on the endomorphisms. Our strategy is to use a classical approach in the study of rational points, that is, answering to a local-global problem. Indeed, for any abelian variety $A/K$, if $\mathfrak{p}$ is a prime of good reduction for $A$, the group $\operatorname{Tors}(A)(K)$ injects into the finite group of $\mathbb{F}_{\mathfrak{p}}$-points of the reduction of $A$ modulo $\mathfrak{p}$. Thus we have that $|\operatorname{Tors}(A)(K)|$ divides $N(\mathfrak{p}):=|\tilde{A}(\mathbb{F}_{\mathfrak{p}})|$, where $\tilde{A}$ is the reduction of $A$ modulo $\mathfrak{p}$. A famous question by Lang asks if the converse is also true. More precisely he asked the following.
@@
Finally, we show how one can apply our main results to study the torsion of modular abelian varieties defined over the rationals, which, by Serre’s conjectures, correspond to weight $2$ newforms. We implement our results in the computational algebra system Magma, in order to produce tentative lists of possible torsion orders for modular abelian abelian varieties of dimension up to $5$, a part of which is presented in Conjectures 4.5 and 4.6. The code and the outputs can be found at [AC26]. Although it is not currently known if the lists are exhaustive, we remark that all the torsion orders appearing in the lists occur for at least one abelian variety, which can be exhibited in terms of the corresponding newform. This data enhances the data that is currently available in LMFDB [18] and in [NIC18]. For example, there exists at least one abelian surface in the isogeny class associated to the level $39$ newform with LMFDB label 39.2.a.b which has torsion order equal to $28$. This abelian variety is not listed in LMFDB and there are no abelian surfaces with torsion order equal to $28$ in the database either.
@@ 2. Setting
A priori, $P_{\mathfrak{p}}^{\lambda}(t)$ has coefficients in $F_{\lambda}$. However, by [RIB76, Theorem 2.1.2], the system $\{\rho_{\lambda}\}_{\lambda}$ is strictly compatible and $F$-rational with exceptional set equal to the set of places of $K$ at which $A$ has bad reduction, so, with possibly these finitely many exceptions, in fact $P_{\mathfrak{p}}^{\lambda}(t)\in F[t]$.
Throughout the paper, for $x\in F_{\lambda}$ (or $F$), we write
$x\equiv 0\pmod{\lambda^{n}}$
whenever the $\lambda$-adic valuation of $x$ is at least $n$.
@@ 3. Main results
Let $A$ be an abelian variety over a number field $K$ of $\text{GL}_{2}$-type. Let $\ell$ be a rational prime and let $\lambda$ be a prime of $F$ over $\ell$. Suppose that, for an integer $n\geq 1$, the congruence
@@
are divisible by the same primes $\ell$ of good reduction for $A$ (i.e. such that, for all places of $K$ above $\ell$, $A$ has good reduction modulo such places). More precisely, for any rational prime $\ell$ of good reduction for $A$, one has
@@
First, let $A^{\prime}$ be an abelian variety that is $K$-isogenous to $A$. By injectivity of the reduction map, which holds for all primes in $\Sigma$ (see [KAT81, Appendix]), $\operatorname{Tors}(A^{\prime})(K)$ injects into the $\mathbb{F}_{\mathfrak{p}}$-points of $A^{\prime}$ for all $\mathfrak{p}\in\Sigma$, thus $|\operatorname{Tors}(A^{\prime})(K)|$ divides $|A^{\prime}(\mathbb{F}_{\mathfrak{p}})|=N(\mathfrak{p})$. Taking the sup over all $A^{\prime}$ that are $K$-isogenous to $A$ and the gcd of all $N(\mathfrak{p})$ as $\mathfrak{p}$ varies in $\Sigma$, we obtain that
$\sup_{A^{\prime}K\text{-isogenous to}A}|\operatorname{Tors}(A^{\prime})(K)|\text{ divides }\gcd_{\mathfrak{p}\in\Sigma}N(\mathfrak{p}).$
Now, let $\ell$ be a prime of good reduction for $A$, dividing all $N(\mathfrak{p})$ for $\mathfrak{p}\in\Sigma$, except those lying above $\ell$. As shown in Equation (1), for every such $\mathfrak{p}$,
@@
The last equality follows from the compatibility of the representations $\{\rho_{\lambda}\}_{\lambda}$, which holds in our setting by [RIB76, Theorem 2.1.2]. Since this quantity is positive, there exists $\lambda_{0}$ lying over $\ell$ with $v_{\lambda_{0}}(P_{\mathfrak{p}}^{\lambda_{0}}(1))=n>0$. By Theorem 3.2, there exists an abelian variety $A^{\prime}$ that is $K$-isogenous to $A$ and such that $v_{\ell}(|\operatorname{Tors}(A^{\prime})(K)|)\geq n\mathfrak{e}_{\lambda_{0}}\mathfrak{f}_{\lambda_{0}}>0$, concluding the proof of the first statement.
@@
So, $v_{\ell}(\gcd N(\mathfrak{p}))\leq n\mathfrak{f}_{\lambda_{0}}\leq v_{\ell}(|\operatorname{Tors}(A^{\prime})(K)|)$. ∎
Remark 3.4.
Notice that the set $\Sigma$ is infinite, since all but finitely many primes are of good reduction and unramified (so in particular $\mathfrak{e}_{p}=1$). However, in order to compute $\gcd_{\mathfrak{p}\in\Sigma}N(\mathfrak{p})$, only finitely many of these primes will be sufficient. See Lemma 4.4 for more details when $K=\mathbb{Q}$.
@@ 4. Modular abelian varieties defined over $\mathbb{Q}$
We now specialise the result of Theorem 3.2 and Corollary 3.3 to simple $\text{GL}_{2}$-type abelian varieties defined over the rational numbers. Let $A$ be an abelian variety of dimension $g$ defined over $\mathbb{Q}$. Let $\operatorname{End}^{0}_{\mathbb{Q}}(A)=\operatorname{End}_{\mathbb{Q}}(A)\otimes_{\mathbb{Z}}\mathbb{Q}$ denote the algebra of endomorphisms of $A$ that are defined over $\mathbb{Q}$, and let $\operatorname{End}_{\overline{\mathbb{Q}}}^{0}(A)=\operatorname{End}_{\overline{\mathbb{Q}}}(A\times_{\mathbb{Q}}\overline{\mathbb{Q}})\otimes_{\mathbb{Z}}\mathbb{Q}$ denote the full endomorphism algebra.
@@
We now investigate the possible geometric behaviours of simple abelian varieties $A$ of $\text{GL}_{2}$-type over $\mathbb{Q}$. As in §2, we write $F\hookrightarrow\operatorname{End}_{\mathbb{Q}}^{0}(A)$, where $F$ is a degree $g$ number field. By [RIB04, Theorem 2.1], $F=\operatorname{End}_{\mathbb{Q}}^{0}(A)$.
@@
Let us assume that $A$ is geometrically simple, so that $\operatorname{End}^{0}_{\overline{\mathbb{Q}}}(A)$ is an Albert algebra. In particular, it is either a division algebra or a field. Let $Z$ be its center. By [RIB04, Proposition 5.1, Proposition 5.2, Theorem 5.3 and Proposition 3.6], $Z$ is totally real, contained in $\operatorname{End}^{0}_{\mathbb{Q}}(A)$, and $\operatorname{End}^{0}_{\overline{\mathbb{Q}}}(A)$ is either a quaternion algebra over $Z$ (PQM case), or it is equal to $Z$, in which case $Z=\operatorname{End}^{0}_{\overline{\mathbb{Q}}}(A)=\operatorname{End}^{0}_{\mathbb{Q}}(A)$ as claimed. ∎
In the non-geometrically simple case, the study of $\operatorname{Tors}(A)$ may be reduced to understanding the torsion of smaller dimensional abelian varieties, albeit defined over larger number fields. In the PQM case, for $A$ of dimension $2$, [LSS+24] classifies all possible torsion subgroups. These first two cases are somewhat special: the generic case is that of $\operatorname{End}^{0}_{\overline{\mathbb{Q}}}(A)=\operatorname{End}^{0}_{\mathbb{Q}}(A)$. Nonetheless, for what follows, we need not distinguish among these cases.
By [RIB04, Theorem 4.4] and Serre’s modularity conjectures [SER87] (proved by Khare–Winterberger, [KW09a, KW09b]), a simple abelian variety $A/\mathbb{Q}$ of $\text{GL}_{2}$-type is modular, hence it is isogenous to the abelian variety $A_{f}$ associated to some newform $f$, with weight $2$ and Nebentypus $\chi$, and coefficient field equal to $F=\operatorname{End}^{0}_{\mathbb{Q}}(A_{f})$. The definition of $A_{f}$ is in [DS05, Definition 6.6.3]). Such an abelian variety has $\lambda$-adic Galois representation isomorphic to that arising from the newform $f$, which we denote by $\rho_{f,\lambda}$. For the construction see [DS05, §9.5]. By [DS05, Theorem 9.6.5], the representations $(\rho_{f,\lambda})$ form a strictly compatible system for all $\lambda$’s, including those of bad reduction for $A_{f}$, with characteristic equation of the Frobenius attached to the rational prime $p$ as follows:
$P_{p}^{\lambda}(t)=1-a_{p}(f)t+\chi(p)pt^{2},$
where $a_{p}(f)$ is the $p$-th Fourier coefficient of $f$.
This discussion allows us to restate Theorem 3.2 and Corollary 3.3 in the following form. The proofs are identical to those in §3, except that we can drop the hypothesis on good reduction of $A_{f}$ at $\ell$, thanks to the compatibility of the $\rho_{f,\lambda}$’s.
@@ 4.1. Implementation and conjectural results
In [AC26], we implement in Magma the results of Theorem 4.2 and Corollary 4.3 to produce a tentative list of torsion orders for simple modular abelian varieties of dimension $g$ over $\mathbb{Q}$. To this end, we access all [^1] For information on the completeness of the data, see https://www.lmfdb.org/ModularForm/GL2/Q/holomorphic/Completeness weight $2$ newforms with level up to $10000$ in LMFDB and list, for each, the following data:
  • the level $N$;
  • the Fourier coefficients $a_{p}$ for $p$ prime up to the Sturm bound (see [STE07, Corollary 9.20]);
  • the Nebentypus $\chi$;
We then use these data to compute, for the non-CM forms (in particular, the associated abelian varieties are non-CM), $P_{p}(1)=P_{p}^{\lambda}(1)=1-a_{p}+p\chi(p)$ for $p$ up to the Sturm bound and such that $\mathfrak{e}_{p}<p-1$, so that torsion injects into the reduction modulo $p$. Recall $P_{p}^{\lambda}(t)$ does not depend on $\lambda$, so we will drop it from the notation. We consider the factorisation in $\mathbb{Z}$ of the norm from $F$ to $\mathbb{Q}$ of $P_{p}(1)$ and, for each rational prime $\ell$ occurring in the factorisation, we compute the sup, for $\lambda$ above $\ell$, of $\ell^{\mathfrak{f}_{\lambda}n}$, where $\mathfrak{f}_{\lambda}$ is the inertia degree of $\lambda$ and $n$ is the $\lambda$-adic valuation of $P_{p}(1)$. Multiplying these terms together, we obtain the predicted torsion order, i.e. the product
 $\prod_{\ell\text{ prime}}\ell^{\sup_{\lambda\mid\ell}(\mathfrak{f}_{\lambda}n)},$
@@
The predicted torsion order is not always equal to the greatest common divisor of the $N(p)$’s, hence Theorem 4.2 and Corollary 4.3 cannot be improved to be exactly equivalent to Katz’s result for elliptic curves. Nonetheless, this is often the case and, using all the data in LMFDB, the two lists computed by the code are exactly the same for abelian varieties of dimension $2$, 33 and $5$. In dimension $4$, the newform with LMFDB label 4830.2.a.ce has an associated abelian variety (up to isogeny) with predicted torsion order equal to $88$, and $\gcd_{p}N(p)=176$, but $88$ does not appear in the list of the $\gcd_{p}N(p)$’s, while $176$ does appear in the list of predicted torsion orders. It is unclear whether, allowing the code to search for higher level newforms, this will remain an exception (and, if so, the only one) or not.
In the following lemma, we prove that using only the Fourier coefficients $a_{p}$ with $p$ up to the Sturm bound ensures the correctness of the predicted torsion orders.
Lemma 4.4.
Let $f$ be a newform of weight $2$, level $N$ and Nebentypus $\chi$. Let st be the Sturm bound for the space of such modular forms. Then, the ideal generated by $a_{p}(f)-(1+\chi(p)p)$ for every $p$ is the same as the ideal generated by the same elements, for $p$ up to st.
Proof.
Let $g$ be an Eisenstein series having $a_{p}(g)=1+\chi(p)p$ for every prime $p$ not dividing the level. This exists, since it can be written as a linear combination of the series in the basis found in [DS05, §4.6] (taking $\varphi=\chi,\psi=1$ and renormalising the coefficients). Let h=fgh=f-g. Then for every prime $p$ not dividing $N$,
$a_{p}(h)=a_{p}(f)-(1+\chi(p)p).$
The claim is equivalent to saying that all the $a_{p}(h)$ are in the ideal generated by the first $a_{p}(h)$ with $p$ up to st. This follows immediately from the Sturm theorem [STE07, Corollary 9.20] applied to the modular form $h$. ∎
We conclude by stating two conjectures, which are suggested by the outcome of the code, on the possible torsion orders of non-CM simple abelian varieties of $\text{GL}_{2}$-type and dimension up to $5$, and on the primes dividing such torsion orders. All the numbers in the table are divisors of the predicted torsion orders computed by the code, and the list of newforms with the respective predicted torsion orders can be found at [AC26]. We do not currently know whether the lists in Conjectures 4.5 and 4.6 are exhaustive. Code and data on how to associate in practice a classical modular form of weight $2$, dimension $2$, and trivial Nebentypus to all genus $2$ curves of $\text{GL}_{2}$-type whose Jacobians are $\mathbb{Q}$-isogenous to the modular abelian varieties associated to $f$ is available at [4].
@@
Let $A/\mathbb{Q}$ be a simple non-CM $g$-dimensional abelian variety of $\text{GL}_{2}$-type. The possible orders of $\operatorname{Tors}(A)(\mathbb{Q})$ are listed below.
@@
Let $A/\mathbb{Q}$ be a simple non-CM $g$-dimensional abelian variety of $\text{GL}_{2}$-type and let $\ell$ be a prime such that $A$ admits a $\ell$-torsion point.
  • If $g=2$, then $\ell\in\{2,3,5,7,11,13,17,19,23,31,37\}$.
  • If $g=3$, then $\ell\in\{2,3,5,7,11,13,17,19,23,29,31,37,83\}$.
  • If $g=4$, then $\ell\in\{2,3,5,7,11,13,17,19,23,29,31,37,41,43,61,71,73\}$.
  • If $g=5$, then $\ell\in\{2,3,5,7,11,13,17,19,23,29,31,41,43,53\}$.